Combining functions. I have to find f over g. f(x)=4x(squared)-19x-63,g(x)=x-7. How do I factor the numerator?
Answer by stanbon(75887)

I have to find f over g. f(x)= 4x(squared)-19x-63, g(x)=x-7. How do I factor the numerator?
---------------
f/g = 4x^2-19x-63 / x-7
f/g = [(4x+9)(x-7)]/[x-7]
f/g = 4x+9
================
Cheers,
Stan H.